# LagCeresa.bib

### v1

@@ 3.1. Motivation and definitions
Notice that finite sets of Lagrangians which are (cylindrically) cobordant are algebraically Lagrangian cobordant, consistently with the comparison between rational and algebraic equivalence. Recall that a planar cobordism always defines a cylindrical cobordism by quotienting $\mathbb{C}$ by a large imaginary translation, therefore without loss of generality we assume there is a cylindrical cobordism $V\subset X\times\mathbb{C}^{*}$ between the finiste sets of Lagrangians $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L^{\prime}_{k^{\prime}})$. Then there is an obvious algebraic Lagrangian cobordism between those finite sets, obtained by gluing two copies of $V\subset X\times\mathbb{C}^{*}$ as shown below.
@@
$\mathcal{C}ob(X)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}algCob(X),$
@@
Given a polarised tropical torus $B$, one can consider the oriented fibered cobordism group $\mathcal{C}ob_{fib}(X(B))$, that is the group generated by fibres $F_{b}$ for $b\in B$ in which relations come from cobordisms all of whose ends are fibres. It follows from flux considerations that there cannot exist an oriented cobordism between $F_{b}$ and $F_{b^{\prime}}$ for $b\neq b^{\prime}$. In fact, Sheridan–Smith prove in [nick&ivan1, Theorem 1.2] that $\mathcal{C}ob_{fib}(X(B))_{hom}$ is divisible (here they consider cylindrical cobordisms). This result mirrors an algebraic result about divisibility of (tropical) Chow groups, namely [bloch_cycles_on_abelian_varieties, Theorem 3.1], of which they prove a tropical version [nick&ivan1, Proposition 3.25]. On the other hand, the diagonal $\Delta\subset X(B)\times X(B)$ provides an obvious oriented algebraic Lagrangian cobordism between any two fibres $F_{b}$ and $F_{b^{\prime}}$, yielding $algCob(X(B))_{fib}\cong\mathbb{Z}$. This is a symplectic manifestation of the fact that any two points on an abelian variety $A$ lie on a curve, hence $\text{Griff}_{0}(A)\cong\mathbb{Z}$. Notice that while certain fluxes do obstruct the existence of algebraic Lagrangian cobordisms, this specific type of flux between fibres does not. This can be understood as a manifestation of Remark 5.11.
As noted in Remark 3.2, in the algebraic setting one can recover an algebraic equivalence of cycles over a curve from an algebraic equivalence of cycles over an abelian variety by considering the flat pullback to a curve between the two relevant points. We end this Section with a Lemma which is an analogous result for algebraic Lagrangian cobordisms:
@@
Because $B$ is polarised, it contains a tropical curve passing through any two given points. For our purposes, we want these curves to admit Lagrangian lifts. We use [nick&ivan1, Lemma 3.28], which states that, given a point $p$ on a polarised tropical torus $B$, there is an open dense subset $U_{p}$ of points in $B$ such that, for any point $q\in U_{p}$, $p$ and $q$ lie on edges of some trivalent tropical curve. Therefore, possibly after perturbing $\Gamma$ by a hamiltonian isotopy, one can take $p$ and $q$ to lie on edges of some trivalent tropical curve $C$ in $B$. The fact that trivalent tropical curves admit Lagrangian lifts follows independently from constructions of Matessi and Mikhalkin in [matessi_lag_pair_of_pants, mikhalkin_trop_to_Lag_corresp], and is also proved in [nick&ivan1, Corollary 4.9] with additional properties of the lift (eg it is graded and locally exact). Let $L_{C}\subset X(B)$ denote such a lift. We now recall a construction of A. Subotic [subotic_monoidal_structure] known as fibrewise sum in a Lagrangian torus fibration with a distinguished Lagrangian section, which in the case of a symplectic polarised torus $X(B)$ is the operation
$\Sigma:\mathcal{F}uk(X(B))\times\mathcal{F}uk(X(B))\longrightarrow\mathcal{F}uk(X(B))$
induced by the correspondence
$\Sigma:=\{(x,p_{1},x,p_{2},x,p_{1}+p_{2}):x\in B,p_{1},p_{2}\in F_{x}\}\subset\overline{X\times X}\times X.$
@@
We now consider the correspondence $\varphi_{13425}(\Delta_{X}\times\Sigma)\subset\overline{X\times X(B)\times X(B)}\times X\times X(B)$, where $\Delta_{X}\subset X\times X$ is the diagonal, and
$\varphi_{13425}$ $:X\times X\times X(B)\times X(B)\times X(B)\longrightarrow X\times X(B)\times X(B)\times X\times X(B)$
$\hskip 56.9055pt(z_{1},z_{2},z_{3},z_{4},z_{5})\hskip 42.67912pt\mapsto\hskip 42.67912pt(z_{1},z_{3},z_{4},z_{2},z_{5})$
@@
is simply the map changing the order of the product components. This correspondence takes Lagrangians in $X\times X(B)\times X(B)$ to Lagrangians in $X\times X(B)$. From this we define $\Gamma^{\prime}$ as $\varphi_{13425}(\Delta_{X}\times\Sigma)(\Gamma\times L_{C})$. We check that $\Gamma^{\prime}(F_{p})=L$, and verifying $\Gamma^{\prime}(F_{q})=L^{\prime}$ is analogous. Notice that, by definition of $\Gamma^{\prime}$,
$\Gamma^{\prime}(F_{p})$ $=\pi_{X}(\Gamma^{\prime}\pitchfork(X\times F_{p}))=\{x\in X:\exists f\in F_{p},(x,f)\in\Gamma^{\prime}))\}$
=$=\{x\in X:\exists f,f_{1},f_{2}\in F_{p},(x,f_{1})\in\Gamma,f_{2}\in L_{C},f=f_{1}+f_{2}\}.$
@@
It remains to notice that because $\Gamma(F_{p})=L$, there exists $f_{1}\in F_{p}$ with $(x,f_{1})\in\Gamma$ if and only if $x\in L$. For $x\in L$, one can take any $f_{2}\in F_{p}\cap L_{C}$ to be the intersection of $F_{p}$ with the zero section in $X(B)$, then $f=f_{1}$ satisfies $(x,f)\in\Gamma^{\prime}$. ∎
3.2. Flux maps from $algCob(X)$
@@
In this Section, we prove that some of the flux maps $\Theta_{\alpha}$ defined in 2 from $\mathcal{C}ob(X)$ (namely those with $\alpha$ of the form $\omega\wedge\beta$) factor through $\mathcal{C}ob(X)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}algCob(X)$.
@@
For any closed form $\beta\in\Omega^{n-3}(X),$ there is a flux-type morphism
$\tilde{\Theta}_{\beta}:$ $\hskip 5.69054ptalgCob(X)_{hom}\longrightarrow\mathbb{R}/\left([\omega^{2}\wedge\beta]\cdot H_{n+1}(X)\right)$
@@
where $\gamma_{0}$ is an $(n+1)$-chain in $X$ with $\partial\gamma_{0}=L$.
@@
Assume there exists a Lagrangian correspondence $\Gamma\subset X\times X(B)$ between the finite sets of Lagrangians $(L_{1},\dots,L_{k})$ and $(L^{\prime}_{1},\dots,L^{\prime}_{k^{\prime}})$. It suffices to show that, if $\gamma_{0}$ is an $(n+1)$-chain in $X$ such that $\partial\gamma_{0}=L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}$, then
$\int_{\gamma_{0}}\omega^{2}\wedge\beta\in[\omega^{2}\wedge\beta]\cdot H_{n+1}(X;\mathbb{Z}).$
We proceed by constructing an $(n+1)$-cycle $C$ in $X\times X(B)$, and show on the one hand that the integral of $(\omega+\omega_{X(B)})^{2}\wedge\pi^{*}_{X}\beta$ on $C$ is the same as an integral over an $(n+1)$-chain $\gamma_{0}$ as above, and on the other hand that $C$ can be continuously deformed to an $(n+1)$-cycle in $X\times F_{p}$. The result will follow. Throughout the proof, we will assume all intersections we consider are transverse, which can always be achieved by a small hamiltonian perturbation of $\Gamma$.
@@
$[0,1]\times(\Gamma\cap(X\times F_{q}))\longrightarrow X\times X(B)$
[…]
where $\overline{P}$ denotes the path $P$ with the orientation reversed, and $x\in X$ while $y\in F_{q}$. Notice that $C_{1}\cup C_{2}$ is an $(n+1)$-chain in $X\times X(B)$ with boundary in $X\times F_{p}$, and more precisely $\pi_{X}(\partial(C_{1}\cup C_{2}))=L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}\subset X$. The fact that the boundary component in $X\times F_{q}$ is empty simply follows from the fact that the boundaries of $\partial C_{1}$ and $\partial C_{2}$ cancel out by construction of $C_{2}$.
Our next step is to construct an $(n+1)$-chain $C_{3}$ inside $X\times F_{p}$ with $\partial C_{3}=\partial(C_{1}\cup C_{2})\subset X\times F_{p}$. Here we use parallel transport again from the fibres $F_{P(t)}$ to $F_{p}$, this time sending $\Gamma\cap(X\times F_{P(t)})$ to $X\times F_{p}$. More explicitly, we take $C_{3}$ to be the image of the map
$C_{1}\hskip 17.07164pt$ $\longrightarrow\hskip 11.38109ptX\times F_{p}$
@@
where $x\in X$, $f\in F_{P(t)}$, and $\tilde{P}_{t}:[0,1]\rightarrow B$ is defined by $\tilde{P}_{t}(s):=P(t(1-s))$. From this construction it follows that $C_{3}$ is $(n+1)$-dimensional except in the trivial case where $C_{1}=C_{2}$ as a set, in which case it is $n$-dimensional, and in particular this would imply $L_{1}+\dots+L_{k}=L^{\prime}_{1}+\dots+L^{\prime}_{k^{\prime}}$. Putting all this together, we have constructed an $(n+1)$-cycle
@@
inside $X\times X(B)$. We now compute the integral
(3.1) $\int_{C}\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta$
@@
$\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)|_{C_{1}}=0$
because $TC_{1}\subset T\Gamma$ and $\Gamma$ is Lagrangian. Although the same doesn’t necessarily hold for $C_{2}$, it is true that
$\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}|_{C_{2}}=0.$
To see this, one can decompose the tangent space to $C_{2}\subset X\times X(B)$ at any point $(x,P(t),f_{P(t)})\in C_{2}$. The component along $X$ lives in an isotropic subspace of $T_{x}X$, because $C_{2}$ is parallel transport of $\Gamma\cap(X\times F_{q})$. By considering the trivialisation
@@
We have now reduced the integral 3.1 to
$\int_{C_{3}}\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta=\int_{C_{3}}\pi_{X}^{*}(\omega^{2}\wedge\beta)=\int_{(\pi_{X})_{*}C_{3}}\omega^{2}\wedge\beta.$
Here the first equality follows from the fact that $C_{3}$ is contained in $X\times F_{p}$ and $F_{p}$ is Lagrangian in $X(B)$. Notice that wa have exhibited an $(n+1)$-chain in $X$, namely $(\pi_{X})_{*}C_{3}$, whose boundary is $L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}$, and over which $\omega^{2}\wedge\beta$ integrated to something which lies in
$[\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta]\cdot H_{n+1}(X\times X(B);\mathbb{Z}).$
The aim of the second computation of 3.1 is to show that the integral over $C$ can actually be reduced to an integral of $\omega^{2}\wedge\beta$ over an $(n+1)$-cycle in $X$. That is, 3.1 actually lies in
$[\omega^{2}\wedge\beta]\cdot H_{n+1}(X;\mathbb{Z})\subset[\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta]\cdot H_{n+1}(X\times X(B);\mathbb{Z}).$
@@
Noticing that $C\subset\pi_{B}^{-1}(P([0,1]))$, this allows us to compute 3.1:
$\int_{C}\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta$ $=\int_{R(1)_{*}C}\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta$
$=\int_{R(1)_{*}C}\pi_{X}^{*}(\omega^{2}\wedge\beta)$
$=\int_{(\pi_{X}\circ R(1))_{*}C}\omega^{2}\wedge\beta.$
It remains to observe that $(\pi_{X}\circ R(1))_{*}C$ is an $(n+1)$-cycle in $X$, therefore we have shown that
$\int_{C}\left(\pi_{X}^{*}\omega+\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{X}^{*}\beta=\int_{(\pi_{X})_{*}C_{3}}\omega^{2}\wedge\beta$
@@
The following is immediate by comparing the definition of the flux maps from $\mathcal{C}ob^{or}(X)$ and $algCob^{or}(X)$.
@@
For $\alpha=\omega\wedge\beta$, the flux morphisms $\Theta_{\alpha}$ and $\tilde{\Theta}_{\beta}$ defined from $\mathcal{C}ob^{or}(X)$ and $algCob^{or}(X)$ factor through the projection map as follows
${\mathcal{C}ob^{or}(X)}$${algCob^{or}(X)}$${\mathbb{R}/\left({[\omega^{2}\wedge\beta]\cdot H_{n+1}(X;\mathbb{Z})}\right).}$$\scriptstyle{\Theta_{\omega\wedge\beta}}$$\scriptstyle{\pi}$$\scriptstyle{\tilde{\Theta}_{\beta}}$

### v2

@@ 4.1. Motivation and definitions
Finite sets of Lagrangians which are (cylindrically) cobordant are algebraically Lagrangian cobordant, consistently with the comparison between rational and algebraic equivalence. Recall that a planar cobordism always defines a cylindrical cobordism by quotienting $\mathbb{C}$ by a large imaginary translation, therefore without loss of generality we assume there is a cylindrical cobordism $V\subset M\times\mathbb{C}^{*}$ between the finite sets of Lagrangians $(L_{1},\dots,L_{k})$ and $(L_{1}^{\prime},\dots,L^{\prime}_{k^{\prime}})$. Then there is an obvious algebraic Lagrangian cobordism between those finite sets, obtained by gluing two copies of $V\subset M\times\mathbb{C}^{*}$ as shown below.
@@
$\mathcal{C}ob(M)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}algCob(M),$
@@
Given a polarised tropical torus $B$, one can consider the oriented fibered cobordism group $\mathcal{C}ob_{fib}(X(B))$, that is the group generated by fibres $F_{b}$ for $b\in B$ in which relations come from cobordisms all of whose ends are fibres. It follows from flux considerations that there cannot exist an oriented cobordism between $F_{b}$ and $F_{b^{\prime}}$ for $b\neq b^{\prime}$. In fact, Sheridan–Smith prove in [nick_ivan1, Theorem 1.2] that $\mathcal{C}ob_{fib}(X(B))_{hom}$ is divisible, which means that for every $L\in\mathcal{C}ob_{fib}(X(B))_{hom}$ and for any $k\in\mathbb{N}$, there exists some $L^{\prime}\in\mathcal{C}ob_{fib}(X(B))_{hom}$ such that $L=k\cdot L^{\prime}$ [^2] Note they work with cylindrical Lagrangian cobordisms. This result mirrors an algebraic result about divisibility of (tropical) Chow groups, namely [bloch_cycles_on_abelian_varieties, Theorem 3.1], of which they prove a tropical version [nick_ivan1, Proposition 3.25]. On the other hand, the diagonal $\Delta\subset\overline{X(B)}\times X(B)$ provides an obvious oriented algebraic Lagrangian cobordism between any two fibres $F_{b}$ and $F_{b^{\prime}}$, yielding $algCob(X(B))_{fib}\cong\mathbb{Z}$. This is a symplectic manifestation of the fact that any two points on an abelian variety $A$ lie on a curve, hence $\text{Griff}_{0}(A)\cong\mathbb{Z}$. Notice that while certain fluxes do obstruct the existence of algebraic Lagrangian cobordisms, this specific type of flux between fibres does not. This is a manifestation of Remark 2.18.
As noted in Remark 4.2, in the algebraic setting one can recover an algebraic equivalence of cycles over a curve from an algebraic equivalence of cycles over an abelian variety by considering the flat pullback to a curve between the two relevant points. We present a Lemma which is an analogous result for algebraic Lagrangian cobordisms:
@@
Because $B$ is polarised, it contains a tropical curve passing through any two given points. For our purposes, we want these curves to admit Lagrangian lifts. We use [nick_ivan1, Lemma 3.28], which states that, given a point $p$ on a polarised tropical torus $B$, there is an open dense subset $U_{p}$ of points in $B$ such that, for any point $q\in U_{p}$, $p$ and $q$ lie on edges of some trivalent tropical curve. Therefore, possibly after perturbing $\Gamma$ by a hamiltonian isotopy, one can take $p$ and $q$ to lie on edges of some trivalent tropical curve $C$ in $B$. The fact that trivalent tropical curves admit Lagrangian lifts follows independently from constructions of Matessi and Mikhalkin in [matessi_lag_pair_of_pants, mikhalkin_trop_to_Lag_corresp], and is also proved in [nick_ivan1, Corollary 4.9] with additional properties of the lift (e.g. it is graded and locally exact). Let $L_{C}\subset X(B)$ denote such a lift. We now recall a construction of A. Subotic [subotic_monoidal_structure] known as fibrewise sum in a Lagrangian torus fibration with a distinguished Lagrangian section. In the case of a symplectic polarised torus $X(B)$, this is the operation of geometric composition with the correspondence
$\Sigma:=\{(b,f_{1},b,f_{2},b,f_{1}+f_{2}):b\in B,f_{1},f_{2}\in F_{b}\}\subset\overline{X(B)\times X(B)}\times X(B).$
@@
We now consider the correspondence $\varphi_{13452}(\Delta_{M}\times\overline{\Sigma})\subset\overline{M}\times X(B)\times X(B)\times\overline{X(B)}\times M$, where $\Delta_{M}\subset\overline{M}\times M$ is the diagonal, and
$\varphi_{13452}$ $:M\times M\times X(B)\times X(B)\times X(B)\longrightarrow M\times X(B)\times X(B)\times M\times X(B)$
$\hskip 56.9055pt(z_{1},z_{2},z_{3},z_{4},z_{5})\hskip 42.67912pt\mapsto\hskip 42.67912pt(z_{1},z_{3},z_{4},z_{5},z_{2})$
@@
is simply the map changing the order of the product components. This correspondence takes Lagrangians in $M\times\overline{X(B)\times X(B)}$ to Lagrangians in $\overline{X(B)}\times M$. From this we define $\Gamma^{\prime}$ as $\varphi_{13452}(\Delta_{M}\times\Sigma)(\Gamma\times L_{C})$.
We check that $\Gamma^{\prime}(F_{p})=L$, and verifying $\Gamma^{\prime}(F_{q})=L^{\prime}$ is analogous. Notice that, by definition of $\Gamma^{\prime}$,
$\Gamma^{\prime}(F_{p})$ $=\pi_{M}(\Gamma^{\prime}\pitchfork(M\times F_{p}))=\{m\in M:\exists f\in F_{p},(m,f)\in\Gamma^{\prime}\}$
$=\{m\in M:\exists f,f_{1},f_{2}\in F_{p},(m,f_{1})\in\Gamma,f_{2}\in L_{C},f=f_{1}+f_{2}\}.$
@@
It remains to notice that because $\Gamma(F_{p})=L$, there exists $f_{1}\in F_{p}$ with $(m,f_{1})\in\Gamma$ if and only if $m\in L$. For $m\in L$, one can take any $f_{2}\in F_{p}\cap L_{C}$ to be the intersection of $F_{p}$ with the zero section in $X(B)$, then $f=f_{1}$ satisfies $(m,f)\in\Gamma^{\prime}$. ∎
4.2. Flux maps from $algCob(M)$
@@
In this Section, we prove that some of the flux maps $\Theta_{\alpha}$ defined in Section 3 from $\mathcal{C}ob(M)$ (namely those with $\alpha$ of the form $\omega\wedge\beta$) factor through $\mathcal{C}ob(M)\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}algCob(M)$. Similarly as for Lagrangian cobordism groups, we denote $algCob^{or}(M)_{hom}$ for the homologically trivial part of the oriented algebraic Lagrangian cobordism group (that is, the kernel of the cycle class map from Proposition 4.4).
@@
For any closed form $\beta\in\Omega^{n-3}(M),$ there is a flux-type morphism
$\tilde{\Theta}_{\beta}:$ $\hskip 5.69054ptalgCob^{or}(M)_{hom}\longrightarrow\mathbb{R}/\left([\omega^{2}\wedge\beta]\cdot H_{n+1}(M)\right)$
@@
where $\gamma_{0}$ is an $(n+1)$-chain in $M$ with $\partial\gamma_{0}=L$.
@@
Assume there exists a Lagrangian correspondence $\Gamma\subset M\times\overline{X(B)}$ between the finite sets of Lagrangians $(L_{1},\dots,L_{k})$ and $(L^{\prime}_{1},\dots,L^{\prime}_{k^{\prime}})$. It suffices to show that, if $\gamma_{0}$ is an $(n+1)$-chain in $M$ with $\partial\gamma_{0}=L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}$, then
$\int_{\gamma_{0}}\omega^{2}\wedge\beta\in[\omega^{2}\wedge\beta]\cdot H_{n+1}(M;\mathbb{Z}).$
We proceed by constructing an $(n+1)$-cycle $C$ in $M\times\overline{X(B)}$, and show on the one hand that the integral of $(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)})^{2}\wedge\pi^{*}_{M}\beta$ on $C$ is the same as an integral over an $(n+1)$-chain $\gamma_{0}$ as above, and on the other hand that $C$ can be continuously deformed to an $(n+1)$-cycle in $M\times F_{p}$. The result will follow. Throughout the proof, we will assume all intersections we consider are transverse, which can always be achieved by a small hamiltonian perturbation of $\Gamma$.
@@
$[0,1]\times(\Gamma\cap(M\times F_{q}))\longrightarrow M\times X(B)$
$\hskip 39.83385pt(t,x,q,y)\hskip 28.45274pt\mapsto\hskip 2.84526pt(x,\varphi_{\overline{P}}(t,y)),$
where $\overline{P}$ denotes the path $P$ with the orientation reversed, and $x\in M$ while $y\in F_{q}$. Notice that $C_{1}\cup C_{2}$ is an $(n+1)$-chain in $M\times X(B)$ with boundary in $M\times F_{p}$, and more precisely $\pi_{M}(\partial(C_{1}\cup C_{2}))=L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}\subset M$. The fact that the boundary component in $M\times F_{q}$ is empty simply follows from the fact that the boundaries of $\partial C_{1}$ and $\partial C_{2}$ cancel out by construction of $C_{2}$.
Our next step is to construct an $(n+1)$-chain $C_{3}$ inside $M\times F_{p}$ with $\partial C_{3}=\partial(C_{1}\cup C_{2})\subset M\times F_{p}$. Here we use parallel transport again from the fibres $F_{P(t)}$ to $F_{p}$, this time sending $\Gamma\cap(M\times F_{P(t)})$ to $M\times F_{p}$. More explicitly, we take $C_{3}$ to be the image of the map
$C_{1}\hskip 17.07164pt$ $\longrightarrow\hskip 11.38109ptM\times F_{p}$
@@
where $x\in M$, $f\in F_{P(t)}$, and $\tilde{P}_{t}:[0,1]\rightarrow B$ is defined by $\tilde{P}_{t}(s):=P(t(1-s))$. From this construction it follows that $C_{3}$ is $(n+1)$-dimensional except in the trivial case where $C_{1}=C_{2}$ as a set, in which case it is $n$-dimensional, and in particular this would imply $L_{1}+\dots+L_{k}=L^{\prime}_{1}+\dots+L^{\prime}_{k^{\prime}}$. Putting all this together, we have constructed an $(n+1)$-cycle
@@
inside $M\times\overline{X(B)}$. We now compute the integral
(4.1) $\int_{C}\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta$
@@
$\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)|_{C_{1}}=0$
because $TC_{1}\subset T\Gamma$ and $\Gamma$ is Lagrangian in $M\times\overline{X(B)}$. Although the same doesn’t necessarily hold for $C_{2}$, it is true that
$\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}|_{C_{2}}=0.$
To see this, one can decompose the tangent space to $C_{2}\subset M\times\overline{X(B)}$ at any point $(x,P(t),f_{P(t)})\in C_{2}$. The component along $M$ lives in an isotropic subspace of $T_{x}M$, because $C_{2}$ is parallel transport of $\Gamma\cap(M\times F_{q})$. By considering the trivialisation
@@
We have now reduced the integral from Equation (4.1) to
$\int_{C_{3}}\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta=\int_{C_{3}}\pi_{M}^{*}(\omega^{2}\wedge\beta)=\int_{(\pi_{M})_{*}C_{3}}\omega^{2}\wedge\beta.$
Here the first equality follows from the fact that $C_{3}$ is contained in $M\times F_{p}$ and $F_{p}$ is Lagrangian in $X(B)$. Notice that wa have exhibited an $(n+1)$-chain in $M$, namely $(\pi_{M})_{*}C_{3}$, whose boundary is $L_{1}+\dots+L_{k}-L^{\prime}_{1}-\dots-L^{\prime}_{k^{\prime}}$, and over which $\omega^{2}\wedge\beta$ integrated to something which lies in
$\left[\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta\right]\cdot H_{n+1}(M\times X(B);\mathbb{Z}).$
The aim of the second computation from Equation (4.1) is to show that the integral over $C$ can actually be reduced to an integral of $\omega^{2}\wedge\beta$ over an $(n+1)$-cycle in $M$. That is, the integral lies in
$\omega^{2}\wedge\beta]\cdot H_{n+1}(M;\mathbb{Z})\subset\left[\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta\right]\cdot H_{n+1}(M\times X(B);\mathbb{Z}).$
@@
Noticing that $C\subset\pi_{B}^{-1}(P([0,1]))$, this allows us to compute the integral in Equation (4.1):
$\int_{C}\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta$ $=\int_{R(1)_{*}C}\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta$
$=\int_{R(1)_{*}C}\pi_{M}^{*}(\omega^{2}\wedge\beta)$
$=\int_{(\pi_{M}\circ R(1))_{*}C}\omega^{2}\wedge\beta.$
It remains to observe that $(\pi_{M}\circ R(1))_{*}C$ is an $(n+1)$-cycle in $M$, therefore we have shown that
$\int_{C}\left(\pi_{M}^{*}\omega-\pi_{X(B)}^{*}\omega_{X(B)}\right)^{2}\wedge\pi_{M}^{*}\beta=\int_{(\pi_{M})_{*}C_{3}}\omega^{2}\wedge\beta$
@@
The following is immediate by comparing the definition of the flux maps from $\mathcal{C}ob^{or}(M)$ and $algCob^{or}(M)$.
@@
For $\alpha=\omega\wedge\beta$, the flux morphisms $\Theta_{\alpha}$ and $\tilde{\Theta}_{\beta}$ defined from $\mathcal{C}ob^{or}(M)$ and $algCob^{or}(M)$ factor through the projection map as follows
$\mathcal{C}ob^{or}(M)$$algCob^{or}(M)$$\mathbb{R}/\left({[\omega^{2}\wedge\beta]\cdot H_{n+1}(M;\mathbb{Z})}\right).$$\scriptstyle{\Theta_{\omega\wedge\beta}}$$\scriptstyle{\pi}$$\scriptstyle{\tilde{\Theta}_{\beta}}$